y = x + 5
4x + y = 20

Answer by stanbon(75887)

y = x + 5
4x + y = 20
------------------
Substitute for "y" in the 2nd equation; then solve for "x":
4x + (x+5) = 20
5x = 15
x = 3
-----------------------
Since y = x+5, y = 3+5, y = 8
---------------------------------
Solution: (3,8)
======================
Cheers,
Stan H.